# 3]Sorbonne Université, CNRS, Université de Paris, Inria, Laboratoire Jacques-Louis Lions UMR 7598, 75005 Paris, France

### v1

@@ 2.4 Phenotype-structured models
The word phenotype refers to the observable features or characteristics of a cell or organism, a term first proposed in [138] to distinguish an individual’s genetic material (the genotype) from its result (the phenotype). Uses of the term have rapidly increased: a naive search within article titles reveals almost 25,000 entries since 2020 (Scopus, 10/10/2024). Used for cell populations, it can refer to physical characteristics (such as shape, size), molecular (such as gene/mRNA/protein expression), and behaviours (such as migration, proliferation). With respect to the historical models above, variations in age (progression through cell cycle), size, and physiological state could all relate to phenotypic structuring.
@@
Growing appreciation into phenotypic variation has led to numerous mathematical models that can be used to understand how it impacts on the spatiotemporal evolution of cellular systems. At a broad level, these attempts can be grouped into two main classes, each with two subclasses:
@@
Discrete population approaches are (usually) predicated on agent-based models (ABMs), where each agent is an individual cell with an assigned phenotypic state; this state is then a determinant of the rules that govern the agent’s behaviour. The flexibility of ABMs, coupled to increased computational power, has led to their rapid expansion in biological modelling (e.g. see Grimm and Railsback 119, Drasdo et al. 71, Metzcar et al. 192, Van Liedekerke et al. 243, Wang et al. 247 and references therein); it is far beyond the scope of this review to cover these in depth, so we just confine to broad remarks. Within each agent the phenotypic state could be discrete or continuous and fixed or variable. Under discrete states, an agent is restricted to occupying one of a finite set of states, whereas under continuous states the phenotype spans a continuum. For varying states, an agent’s phenotypic state can change over time, e.g. as a result of interactions with other cells or the environment. For example, this could be enacted through equipping each cell with a dynamical system to describe intracellular signalling; various open source toolkits are available for implementing such ABMs – e.g. see the review by [192].
@@
Continuous approaches assume that the population can be represented by a continuum, i.e. as a density distribution. As above, these can be further divided into those with discrete or continuous phenotype states. Under the former, the overall dynamics are represented by a system of (typically coupled) evolution equations, each ODE or PDE describing the evolution of the density of individuals with a certain phenotype. Models of this type have become popular in recent years, in particular to describe how cellular heterogeneity can impact on cancer progression, invasion, and treatment (e.g. Gatenby and Gawlinski 110, Fedotov and Iomin 94, Gerlee and Nelander 117, Pham et al. 201, Stepien et al. 224, Strobl et al. 220, Crossley et al. 66). Relying on just a few phenotype states – such as a binary description, see Figure 2(d) – these models fall into the general class of low(ish) order ODE/PDE systems and can be approached analytically with standard techniques (linear stability, bifurcation analyses etc). Continuous models in which the phenotype enters as a continuous structuring variable represent a further step up in model complexity, but also benefit from a capacity to capture the complex heterogeneity often present in a population. Many models of this type fall into the PS-PIDE structure (2), and we will be using the rest of this review to describe these in more detail.
3 PDEs to PS-PIDEs for biological movement
@@
The PDE (1) underpins many models and has been tailored to describe various types of movement. We use this section to discuss the extension of some classic PDE models of form (1) into the PS-PIDE form (2). In cell systems, dynamics are regulated by the interactions between cells and other components of the extracellular environment. For this reason, the dependency on $({t},{\bm{x}})$ of the terms ${\bm{A}_{\bm{x}}}$, ${D_{\bm{x}}}$, ${\bm{A}_{\bm{y}}}$, ${D_{\bm{y}}}$, and $F$ in (2) may be mediated by macroscopic quantities, in addition to the density $\rho$, modelling other biotic (e.g. densities of cells of other populations) and abiotic factors (e.g. concentrations of soluble molecules and densities of insoluble polymers) in the microenvironment. These in turn may satisfy their own evolution equations, but for the sake of conciseness we will not discuss them here. The definitions of ${\bm{A}_{\bm{x}}}$ and ${D_{\bm{x}}}$ will depend on the type of movement one seeks to represent, and in the following we will consider some standard forms corresponding to different cell migration models. Note also that while in this section we retain the notation introduced in Section 1, considering a $p$-dimensional phenotypic structure ($p\geq 1$), the majority of the modelling works in the extant literature cited below consider a one-dimensional phenotypic structure ($p=1$). As a further note, the PDEs and PS-PIDEs presented in this section – when posed on bounded domains – are subject to zero-flux boundary conditions, unless stated otherwise.
@@ 3.1 Diffusion-based movement models
Figure 3: Illustration of travelling waves from the Fisher-KPP model. Illustrative outputs of the Fisher-KPP model (5)-(6), with ${\bm{x}}\equiv x\in[0,50]$, subject to zero-flux boundary conditions. In particular, evolution to travelling waves is shown for the rescaled PDE (28), complemented with an initial condition of form (29), with (a) $\varepsilon=1$, (b) $\varepsilon=0.1$, (c) $\varepsilon=0.01$. In each case the density $\rho(t,x)$ is plotted at times $t=5$ (blue lines), $t=10$ (green lines), $t=15$ (red lines), and $t=20$ (red lines). We see an evolution of the solutions towards a travelling wave profile with wave speed $c=2$ (numerically calculated wave speeds $v$ reported for each simulation at times $t=5$ and $t=20$). As $\varepsilon\rightarrow 0$, this occurs on a fast timescale and the profile becomes increasingly step-like, from $\rho\equiv 1$ across the wave to $\rho\equiv 0$ ahead of the wave, in agreement with analytical predictions.
Linear (or Fickian) diffusion describes the effect of Brownian motion at the macroscopic scale (i.e. the undirected motion of particles randomly moving in space), and follows from Fick’s first law [95]. The relative simplicity of this form has led to its adoption in many models of biological motion – both of cells and animals – and is a component of a mathematical biology mainstay: the Fisher-KPP equation, which we adopt as an exemplar.
Proposed in [97] and [146], the Fisher-KPP equation forms travelling waves for a population that expands about its focal point (see Figure 3) and lies at the basis of models for processes that range from ecological invasions to tumour growth. In its textbook statement, linear diffusion is combined with logistic growth so that one defines $D_{\bm{x}}\equiv D$, ${\bm{A}_{\bm{x}}}\equiv 0$, $F:=\rho R(\rho)$ in (1) and obtains
@@
where $r\in\mathbb{R}^{+}_{*}$ is the intrinsic growth rate (i.e. the growth rate when $\rho=0$) and $k\in\mathbb{R}^{+}_{*}$ [^10] In the remainder of the paper, we will use the notation +:=+$\mathbb{R}^{+}_{*}:=\mathbb{R}^{+}\setminus\{0\}$, where $\mathbb{R}^{+}$ is the set of non-negative real numbers. is the local carrying capacity.
@@
A conventional reading of (5) suggests a single (i.e. homogeneous) population model, but the original works of [97] and [146] were, in fact, for structured (i.e. heterogeneous) populations. For example, Fisher was investigating [^11] In commenting on this background to (5), we should mention the roots to eugenics: [97] appeared in the (discontinued) Annals of Eugenics and Fisher was an editor of the journal. This field became increasingly politicised and notorious across the first decades of the 20th century, so our discussion here is just to highlight that the common interpretation of (5) as a single population model had, in its origin, a structured population in mind. the spread of advantage-gaining mutations. Under this motivation, the function $\rho$ in (5) describes the density of the ‘daughter’ population of an allelomorph pair, with the density of the ‘parent’ population set by $k-\rho$ under an assumption that the combined daughter and parent density is constant. The travelling waves that can arise (see Figure 3) then describe how the daughter population expands to replace the parent. As such, the well-known Fisher-KPP equation – as originally intended – described the evolving heterogeneous structure in a population, albeit under a highly simple scenario.
@@
Figure 4: Illustration of travelling waves from a PS-PIDE model of diffusion-based movement. Illustrative output of the PS-PIDE model of diffusion-based movement (7). Here, we consider one dimension in each of physical and phenotype space, such that ${\bm{x}}\equiv x\in[0,L]$ with $L\in\mathbb{R}^{+}_{*}$ and ${\bm{y}}\equiv y\in[0,1]$. Zero-flux conditions are imposed at the boundaries and the population is initially set such that $n(0,x,y)=0.1e^{-100x}$. We take $D(y):=y$ and define $R(y,\rho):=r(y)-\kappa\rho$, with $r(y):=10(1-y)$ and $\kappa=10$. These definitions describe a trade-off in which the most motile population members ($y\approx 1$) have the lowest proliferation rate, while the least motile members ($y\approx 0$) have the highest proliferation rate. Simulations are for (a) $\bar{D}=10^{-1}$, (b) $\bar{D}=10^{-3}$, and (c) $\bar{D}=10^{-5}$. Note that $L=100$ in (a-b) and $L=200$ in (c), due to the longer time it takes to evolve to the travelling wave profile in (c). Top row displays the phenotype density function $n(t,x,y)$ with the different colour-themes corresponding to different times (colour bars in (c)). Central row displays the density $\rho(t,x)$ at the same times, showing the evolution to travelling wave solutions. Bottom row displays the density $\rho$ at (a-b) $t=25$ and (c) $t=50$, where the colour code indicates the locally prevailing phenotype, $\bar{\bm{y}}(t,{\bm{x}})\equiv\bar{y}(t,x)$, across the wave for (a-b) $t=25$ and (c) $t=50$. Broadly speaking, lowering the diffusivity in phenotype space (i.e. reducing the value of the rate of phenotypic changes $\bar{D}$) leads to an increasingly structured profile, where more mobile members dominate the front of the wave and more proliferative members dominate the rear.
@@
How can we generalise this to incorporate even greater structuring, i.e. a population continuously structured across some phenotypic space? Let us consider the phenotype density $n(t,{\bm{x}},{\bm{y}})$ and allow transitions to take place between the phenotypic states ${\bm{y}}\in{\cal{Y}}$. Taking the PS-PIDE form (2), it is common to assume that phenotypic changes can be described by a linear diffusion process, so that ${\bm{A}}_{\bm{y}}\equiv 0$ and $D_{\bm{y}}\equiv\bar{D}$. In cell populations, such a modelling approach is natural, say, if changes are driven by stochastic fluctuations of signalling pathways [60]. If the phenotype of population members impact on both their motility and proliferation rate, a natural extension of (5) is then to the following PS-PIDE formulation:
@@
In the framework of PS-PIDE models of form (7), the parameter $\bar{D}\in\mathbb{R}^{+}$ can be regarded as the rate of phenotypic changes. Furthermore, the function $R({\bm{y}},\rho)$ models the net proliferation rate (i.e. the difference between the rate of proliferation and the rate of death) of population members in the phenotypic state ${\bm{y}}$ under the local environmental conditions corresponding to the density $\rho$. Hence, it can be regarded as the fitness function or the fitness landscape of the population at position ${\bm{x}}$ [45].
Depending on the assumptions that are made on the functions and parameters of the model, PS-PIDEs of form (7) can produce phenotype-structured travelling waves, wherein the locally prevailing phenotype, $\bar{\bm{y}}$, at different positions – which we define as the point $\bar{\bm{y}}(t,{\bm{x}})$ such that ${n(t,{\bm{x}},\bar{\bm{y}}(t,{\bm{x}}))=\max_{{\bm{y}}\in\mathcal{Y}}n(t,{\bm{x}},{\bm{y}})}$ – varies across the wave (see Figure 4). These can be seen as phenotypically heterogeneous waves of invasion, whereby population members in different phenotypic states dominate different parts of the wave.
Variants of PS-PIDE models of form (7) for situations in which the phenotype structure is not linked to the motility of the population members (i.e. when $D({\bm{y}})\equiv D$) have received attention from the mathematical community for the richness of the properties of their solutions [7, 157]. On the other hand, variants of (7) for scenarios where the phenotype structure is linked to the motility of the population members but not to the rates at which they proliferate, i.e. PS-PIDE models of form
@@
was proposed in [29]. More recently, related models that take into account the effect of spatial heterogeneity in the surrounding environment on the proliferation rate [163] and motility [252] of the population members have also been considered. Over the last decade, the PS-PIDE model (8)-(9) has received considerable attention, both from a modelling and an analytical point of view. In particular, from the modelling perspective, it has provided a robust mechanistic explanation for spatial sorting observed in cane toad invasion: highly motile individuals are found to reside at the edge of the invasion front [29, 38, 28, 26, 233]. From the analytical standpoint, it has become a prototypical example of PS-PIDE models that admit accelerating-front solutions (i.e. travelling-front solutions that accelerate over time) [38, 28, 35].
Moreover, when spatial dynamics are ignored, i.e. $n(t,{\bm{x}},{\bm{y}})\equiv n(t,{\bm{y}})$, PS-PIDE models of form (7) reduce to the following well-investigated class of non-local PDE models of evolutionary dynamis  [^13] It is common to refer to non-local PDE models of form (10) as mutation-selection models (e.g. see Lorenzi and Pouchol 171) or replicator-mutator equations (e.g. see Alfaro and Veruete 18) or Lotka-Volterra parabolic equations (e.g. see Perthame and Barles 200). On the other hand, when $R({\bm{y}},\rho)\equiv R(\rho)$ with $R(\rho)$ of form (6) or of a related form, non-local PDE models of this type are usually referred to as non-local Fisher-KPP equation (e.g. see Berestycki et al. 39), since they can be regarded as a non-local version of the Fisher-KPP model (5)-(6). However, note that in the literature the term non-local Fisher-KPP equation is sometimes used, by extension, to refer to PS-PIDE models of form (10) as well.
@@
where now $\rho(t)$ is the (total) population size at time $t$. Models of this type have been widely used in theoretical studies into the evolutionary dynamics of phenotype-structured populations, for different biological contexts [202]. Focusing on cell populations, a possible form of the fitness function $R({\bm{y}},\rho)$ that has been employed, e.g. in [1, 155], to investigate the evolutionary dynamics of cancer cells in spatially homogeneous scenarios is
@@
Here the point ${\bm{\varphi}}\in\mathcal{Y}$ models the phenotypic state corresponding to the fitness peak (i.e. the fittest phenotype) and the parameter $\gamma\in\mathbb{R}^{+}_{*}$ is linked to the maximum fitness (i.e. the fitness of the fittest phenotype). Moreover, the saturating term $-\kappa\rho$ in (11) models the limitations on population growth imposed by carrying capacity constraints, and the parameter $\kappa\in\mathbb{R}^{+}_{*}$ is inversely related to the carrying capacity of the population. Of course, when modelling the evolutionary dynamics of cell populations, alternatives to (11) can also be considered for the fitness function, as done, for instance, in [165, 62, 112].
@@
As illustrated by the plots in Figure 5, when $p=1$, and thus ${\bm{y}}\equiv y$ and ${\bm{\varphi}}\equiv\varphi$, non-local PDE models of form (10) complemented with definitions (11)-(12) admit Gaussian-like solutions with a variance proportional to $\bar{D}$ and a mean, $\bar{y}(t)$, which represents the mean or prevailing phenotype in the population at time $t$, that converges to $\varphi$ as $t\to\infty$ [1, 9, 61]. This provides a possible mathematical formalisation of two key facts in evolutionary biology: the rate at which individuals undergo phenotypic changes, $\bar{D}$, impacts on the phenotypic variance (i.e. the level of phenotypic heterogeneity) of the population; and the mean or prevailing phenotype in the population will ultimately be the fittest phenotype $\varphi$. Moreover, as discussed in detail in Section 4.2.2 and as illustrated again by Figure 5, appropriately rescaled versions of these models can exhibit concentration phenomena – i.e. their solutions can become concentrated as weighted infinitely sharp Gaussians, that is, weighted Dirac masses – formalising in mathematical terms the idea that the population shows only one trait, and thus remains monomorphic over time [75, 202]. We note that other forms of explicit and semi-explicit solutions to variants of the non-local PDE (10) have also been obtained (e.g. see Alfaro and Carles 3, Alfaro and Carles 4) and cases where there are multiple fitness peaks have also been analytically investigated (e.g. see Alfaro and Veruete 18, Lorenzi and Pouchol 171).
@@
Figure 5: Illustration of Gaussian-like solutions and concentration phenomena from a non-local PDE model of evolutionary dynamics. Illustrative output of the non-local PDE model of evolutionary dynamics (10)-(12), with ${\bm{y}}\equiv y\in[-5,5]$, subject to zero-flux boundary conditions. In particular, Gaussian-like solutions and concentration phenomena are shown for the rescaled non-local PDE (35), under definitions (11)-(12) (with $\kappa=0.1$, $\gamma=10$, and $\varphi=1$), subject to initial condition (37) (with $\rho^{0}=60$ and $\bar{y}^{0}=-1$). Simulations are for (a) $\varepsilon=1$, (b) $\varepsilon=0.1$, and (c) $\varepsilon=0.01$. Top panels display the dynamics of the rescaled phenotype density $n(t,y)/\rho(t)$; bottom panels display the phenotype density $n(t,y)$ (red solid line) at select times. We compare the numerical solution with the analytically predicted one (black dashed lines) – i.e. the Gaussian-like solution (38) with mean $\bar{y}_{\varepsilon}(t)$ and weight $\rho_{\varepsilon}(t)$ obtained by solving numerically the Cauchy problem (39). We see the solution to be of a Gaussian form, which becomes concentrated as a weighted increasingly sharp Gaussian as $\varepsilon$ gets smaller, corroborating the emergence of concentration phenomena of type (40). We see also that the mean or prevailing phenotype $\bar{y}(t)$ converges to $y=\varphi$ (black dashed lines, top panels) as $t\to\infty$, corresponding to selection of the fittest trait.
@@
Finally, for a spatially heterogeneous environment – for example a population surrounded by non-uniformly distributed abiotic factors (nutrients, therapeutic agents, etc for cell populations) that impact on proliferation and death rates – a natural extension of PS-PIDE models of form (7) is

### v2

@@ 2.4 Phenotype-structured models
The word phenotype refers to the observable features or characteristics of a cell or organism, a term first proposed by [141] to distinguish an individual’s genetic material (the genotype) from its result (the phenotype). Used for cell populations it can refer to physical characteristics (such as shape, size), molecular (such as gene/mRNA/protein expression), and behaviours (such as migration, proliferation). With respect to the historical models above, variations in age (progression through cell cycle), size, and physiological state could all relate to phenotypic structuring.
@@
Growing appreciation for phenotypic variation has resulted in numerous mathematical models that can be used to understand the impact it has on the spatiotemporal evolution of cellular systems. At a broad level, these attempts can be grouped into two main classes, each with two subclasses:
@@
Discrete population approaches are (usually) predicated on agent-based models (ABMs), where each agent is an individual cell with an assigned phenotypic state; this state is then a determinant of the rules that govern the agent’s behaviour. The flexibility of ABMs, coupled to increased computational power, has led to their rapid expansion in biological modelling (e.g. see Grimm and Railsback 117, Drasdo et al 82, Metzcar et al 189, Van Liedekerke et al 241, Wang et al 249 and references therein); it is far beyond the scope of this review to cover these in depth, so we just confine to broad remarks. Within each agent the phenotypic state could be discrete or continuous and fixed or variable. Under discrete states, an agent is restricted to occupying one of a finite set of states, whereas under continuous states the phenotype spans a continuum. For varying states, an agent’s phenotypic state can change over time, e.g. as a result of interactions with other cells or the environment. For example, this could be enacted through equipping each cell with a dynamical system to describe intracellular signalling; various open source toolkits are available for implementing such ABMs – e.g. see the review by [189].
@@
Continuous approaches assume that the population can be represented by a continuum, i.e. as a density distribution. As above, these can be further divided into those with discrete or continuous phenotype states. Under the former, the overall dynamics are represented by a system of (typically coupled) evolution equations, each ODE or PDE describing the evolution of the density of individuals with a certain phenotype. Models of this type have become popular in recent years, in particular to describe how cellular heterogeneity can have an impact on cancer progression, invasion, and treatment (e.g. Gatenby and Gawlinski 114, Fedotov and Iomin 93, Gerlee and Nelander 115, Pham et al 211, Stepien et al 228, Strobl et al 229, Crossley et al 68). Relying on just a few phenotype states – such as a binary description, see Figure 2(d) – these models fall into the general class of low(ish) order ODE/PDE systems and can be approached analytically with standard techniques (linear stability, bifurcation analyses etc). Continuous models in which the phenotype enters as a continuous structuring variable represent a further step up in model complexity, but also benefit from a capacity to capture the complex heterogeneity often present in a population. Many models of this type fall into the PS-PDE structure (2), and we will be using the rest of this review to describe these in more detail.
3 PDEs to PS-PDEs for biological movement
@@
The PDE (1) underpins many models and has been tailored to describe various types of movement. We use this section to discuss the extension of some classic PDE models of form (1) into the PS-PDE form (2). In cell systems, dynamics are regulated by the interactions between cells and other components of the extracellular environment. For this reason, the dependency on $({t},{\bm{x}})$ of the terms ${\bm{A}_{\bm{x}}}$, ${D_{\bm{x}}}$, ${\bm{A}_{\bm{y}}}$, ${D_{\bm{y}}}$, and $F$ in (2) may be mediated by macroscopic quantities, in addition to the density $\rho$, modelling other biotic (e.g. densities of cells of other populations) and abiotic (e.g. concentrations of soluble molecules and densities of insoluble polymers) factors in the microenvironment. These in turn may satisfy their own evolution equations, but for the sake of conciseness we will not discuss them here. The definitions of ${\bm{A}_{\bm{x}}}$ and ${D_{\bm{x}}}$ will depend on the type of movement one seeks to represent, and in the following we will consider some standard forms corresponding to different cell migration models. Note also that while in this section we retain the notation introduced in Section 1, considering a $p$-dimensional phenotypic structure ($p\geq 1$), the majority of the modelling works in the extant literature cited below consider a one-dimensional phenotypic structure ($p=1$). As a further note, the PDEs and PS-PDEs presented in this section – when posed on bounded domains – are subject to zero-flux boundary conditions, unless stated otherwise.
@@ 3.1 Diffusion-based movement models
Figure 3: Illustration of travelling wave like solutions from the Fisher-KPP model. Illustrative outputs of the Fisher-KPP model (5)-(6). In particular, evolution towards a travelling wave form is shown for the rescaled PDE (28), complemented with an initial condition of form (29), with (a) $\varepsilon=1$, (b) $\varepsilon=0.1$, (c) $\varepsilon=0.01$. In each case the density $\rho(t,x)$ is plotted at times $t=5$ (blue lines), $t=10$ (green lines), $t=15$ (red lines), and $t=20$ (red lines). We see an evolution of the solutions towards a travelling wave profile with wave speed $c=2$ (numerically calculated wave speeds $v$ reported for each simulation at times $t=5$ and $t=20$). As $\varepsilon\rightarrow 0$, this occurs on a fast timescale and the profile becomes increasingly step-like, from $\rho\equiv 1$ across the wave to $\rho\equiv 0$ ahead of the wave, in agreement with analytical predictions. While travelling waves are usually considered on an unbounded domain, the numerical method requires a bounded domain. For these simulations we have set ${\bm{x}}\equiv x\in[0,50]$, subject to zero-flux boundary conditions.
Linear (or Fickian) diffusion describes the effect of Brownian motion at the macroscopic scale (i.e. the undirected motion of particles randomly moving in space), and follows from Fick’s first law [96]. The relative simplicity of this form has led to its adoption in many models of biological motion and is a component of a mathematical biology mainstay: the Fisher-KPP equation, which we adopt as an exemplar.
Proposed by [98] and [153], the Fisher-KPP equation forms travelling waves for a population that expands about its focal point (see Figure 3), and lies at the basis of models for processes that range from ecological invasions to tumour growth. In its textbook statement, linear diffusion is combined with logistic growth so that one defines $D_{\bm{x}}\equiv D$, ${\bm{A}_{\bm{x}}}\equiv 0$, $F:=\rho R(\rho)$ in (1) and obtains
@@
where $r\in\mathbb{R}^{+}_{*}$ [^11] In the remainder of the paper, we will use the notation +:=+$\mathbb{R}^{+}_{*}:=\mathbb{R}^{+}\setminus\{0\}$, where $\mathbb{R}^{+}$ is the set of non-negative real numbers. is the intrinsic growth rate (i.e. the growth rate when $\rho=0$) and $k\in\mathbb{R}^{+}_{*}$ is the local carrying capacity.
@@
A conventional reading of (5) suggests a single (i.e. homogeneous) population model, but the original works of [98] and [153] were, in fact, for structured (i.e. heterogeneous) populations. For example, Fisher was investigating [^12] In commenting on this background to (5), we should mention the roots to eugenics: [98] appeared in the (discontinued) Annals of Eugenics and Fisher was an editor of the journal. This field became increasingly politicised and notorious across the first decades of the 20th century, so our discussion here is just to highlight that the common interpretation of (5) as a single population model had, in its origin, a structured population in mind. the spread of advantage-gaining mutations. Under this motivation, the function $\rho$ in (5) describes the density of the ‘daughter’ population of an allelomorph pair, with the density of the ‘parent’ population set by $k-\rho$ under an assumption that the combined daughter and parent density is constant. The travelling waves that can arise (see Figure 3) then describe how the daughter population expands to replace the parent. As such, the well-known Fisher-KPP equation – as originally intended – described the evolving heterogeneous structure in a population, albeit under a highly simplified scenario.
@@
Figure 4: Illustration of travelling wave like solutions from a PS-PDE model of diffusion-based movement. Illustrative output of the PS-PDE model of diffusion-based movement (7). Here, we consider one dimension in each of physical and phenotype space, such that ${\bm{x}}\equiv x\in[0,L]$ with $L\in\mathbb{R}^{+}_{*}$ and ${\bm{y}}\equiv y\in[0,1]$. Zero-flux conditions are imposed at the boundaries and the population is initially set such that $n(0,x,y)=0.1e^{-100x}$. We take $D(y):=y$ and define $R(y,\rho):=r(y)-\kappa\rho$, with $r(y):=10(1-y)$ and $\kappa=10$. These definitions describe a trade-off in which the most motile population members ($y\approx 1$) have the lowest proliferation rate, while the least motile members ($y\approx 0$) have the highest proliferation rate. Simulations are for (a) $\bar{D}=10^{-1}$, (b) $\bar{D}=10^{-3}$, and (c) $\bar{D}=10^{-5}$. Note that $L=100$ in (a-b) and $L=200$ in (c), due to the longer time it takes to evolve to the travelling wave profile in (c). Top row displays the phenotype density function $n(t,x,y)$ with the different colour-themes corresponding to different times (colour bars in (c)). Central row displays the density $\rho(t,x)$ at the same times, showing the evolution to travelling wave solutions. Bottom row displays the density $\rho$ at (a-b) $t=25$ and (c) $t=50$, where the colour code indicates the locally prevailing phenotype, $\bar{\bm{y}}(t,{\bm{x}})\equiv\bar{y}(t,x)$, across the wave for (a-b) $t=25$ and (c) $t=50$. Broadly speaking, lowering the diffusivity in phenotype space (i.e. reducing the value of the rate of phenotypic changes $\bar{D}$) leads to an increasingly structured profile, where more mobile members dominate the front of the wave and more proliferative members dominate the rear.
@@
How can we generalise this to incorporate even greater structuring, i.e. a population continuously structured across some phenotypic space? Let us consider the phenotype density $n(t,{\bm{x}},{\bm{y}})$ and allow transitions to take place between the phenotypic states ${\bm{y}}\in{\cal{Y}}$. Taking the PS-PDE form (2), it is common to assume that phenotypic changes can be described by a linear diffusion process, so that ${\bm{A}}_{\bm{y}}\equiv 0$ and $D_{\bm{y}}\equiv\bar{D}$. In cell populations, such a modelling approach is natural, say, if changes are driven by stochastic fluctuations of signalling pathways [60]. If the phenotype of population members affects both their motility and proliferation rate, a natural extension of (5) is then to the following PS-PDE formulation:
@@
In the framework of PS-PDE models of form (7), the parameter $\bar{D}\in\mathbb{R}^{+}$ can be regarded as the rate of phenotypic changes. Furthermore, the function $R({\bm{y}},\rho)$ models the net proliferation rate (i.e. the difference between the rate of proliferation and the rate of death) of population members in the phenotypic state ${\bm{y}}$ under the local environmental conditions corresponding to the density $\rho$. Hence, it can be regarded as the fitness function or the fitness landscape of the population [42].
Depending on the assumptions that are made on the functions and parameters of the model, PS-PDEs of form (7) can produce phenotype-structured travelling waves, wherein the locally prevailing phenotype, $\bar{\bm{y}}$, at different positions – which we define as the point $\bar{\bm{y}}(t,{\bm{x}})$ such that ${n(t,{\bm{x}},\bar{\bm{y}}(t,{\bm{x}}))=\max_{{\bm{y}}\in\mathcal{Y}}n(t,{\bm{x}},{\bm{y}})}$ – varies across the wave (see Figure 4). These can be seen as phenotypically heterogeneous waves of invasion, whereby population members in different phenotypic states dominate different parts of the wave.
Variants of PS-PDE models of form (7) for situations in which the phenotype structure is not linked to the motility of the population members (i.e. when $D({\bm{y}})\equiv D$) have received attention from the mathematical community for the richness of the properties of their solutions [16, 160]. On the other hand, variants of (7) for scenarios where the phenotype structure is linked to the motility of the population members but not to the rates at which they proliferate/die, i.e. PS-PDE models of form
@@
was proposed by [26]. More recently, related models that take into account the effect of spatial heterogeneity in the surrounding environment on the proliferation rate [156] and motility [254] of the population members have also been considered. Over the last decade, the PS-PDE model (8)-(9) has received considerable attention, both from a modelling and an analytical point of view. In particular, from the modelling perspective, it has provided a robust mechanistic explanation for spatial sorting observed in cane toad invasion: highly motile individuals are found to reside at the edge of the invasion front [26, 29, 34, 32, 238]. From the analytical standpoint, it has become a prototypical example of PS-PDE models that admit accelerating-front solutions (i.e. travelling-front solutions that accelerate over time) [29, 34, 35].
Moreover, when spatial dynamics are ignored, i.e. $n(t,{\bm{x}},{\bm{y}})\equiv n(t,{\bm{y}})$, PS-PDE models of form (7) reduce to the following well-investigated class of non-local PDE models of evolutionary dynamis  [^14] It is common to refer to non-local PDE models of form (10) as mutation-selection models, e.g. see [166], or replicator-mutator equations, e.g. see [5], or Lotka-Volterra parabolic equations, e.g. see [208]. On the other hand, when $R({\bm{y}},\rho)\equiv R(\rho)$ with $R(\rho)$ of form (6) or of a related form, non-local PDE models of this type are usually referred to as non-local Fisher-KPP equation, e.g. see [28], since they can be regarded as a non-local version of the Fisher-KPP model (5)-(6). However, note that in the literature the term non-local Fisher-KPP equation is sometimes used, by extension, to refer to non-local PDE models of form (10) as well.
@@
where now $\rho(t)$ is the (total) population size at time $t$. Models of this type have been widely used in theoretical studies into the evolutionary dynamics of phenotype-structured populations, for different biological contexts [205]. Focusing on cell populations, a possible form of the fitness function $R({\bm{y}},\rho)$ that has been employed, e.g. by [9, 168], to investigate the evolutionary dynamics of cancer cells in spatially homogeneous scenarios is
@@
Here the point ${\bm{\varphi}}\in\mathcal{Y}$ models the phenotypic state corresponding to the fitness peak (i.e. the fittest phenotype) and the parameter $\gamma\in\mathbb{R}^{+}_{*}$ is linked to the maximum fitness (i.e. the fitness of the fittest phenotype). Moreover, the saturating term $-\kappa\rho$ in (11) models the limitations on population growth imposed by carrying capacity constraints, and the parameter $\kappa\in\mathbb{R}^{+}_{*}$ is inversely related to the carrying capacity of the population. Of course, when modelling the evolutionary dynamics of cell populations, alternatives to (11) can also be considered for the fitness function, as done, for instance, by [175, 59, 118].
@@
As illustrated by the plots in Figure 5, when $p=1$, and thus ${\bm{y}}\equiv y$ and ${\bm{\varphi}}\equiv\varphi$, non-local PDE models of form (10) complemented with definitions (11),(12) admit Gaussian-like solutions with a variance proportional to $\bar{D}$ and a mean, $\bar{y}(t)$, which represents the mean or prevailing phenotype in the population at time $t$, that converges to $\varphi$ as $t\to\infty$ [9, 15, 61]. This provides a possible mathematical formalisation of two key facts in evolutionary biology: the rate at which individuals undergo phenotypic changes, $\bar{D}$, impacts on the phenotypic variance (i.e. the level of phenotypic heterogeneity) of the population; and the mean or prevailing phenotype in the population will ultimately be the fittest phenotype $\varphi$. Moreover, as discussed in detail in Section 4.2.2 and as illustrated again by Figure 5, appropriately rescaled versions of these models can exhibit concentration phenomena – i.e. their solutions can become concentrated as weighted infinitely sharp Gaussians, that is, weighted Dirac masses – formalising in mathematical terms the idea that the population shows only one trait, and thus remains monomorphic over time [78, 205]. We note that other forms of explicit and semi-explicit solutions to variants of the non-local PDE (10) have also been obtained, e.g. see [3, 4], and cases where there are multiple fitness peaks have also been analytically investigated, e.g. see [5, 166].
@@
Figure 5: Illustration of Gaussian-like solutions and concentration phenomena from a non-local PDE model of evolutionary dynamics. Illustrative output of the non-local PDE model of evolutionary dynamics (10)-(12), with ${\bm{y}}\equiv y\in[-5,5]$, subject to zero-flux boundary conditions. In particular, Gaussian-like solutions and concentration phenomena are shown for the rescaled non-local PDE (35), under definitions (11),(12) (with $\kappa=0.1$, $\gamma=10$, and $\varphi=1$), subject to initial condition (37) (with $\rho^{0}=60$ and $\bar{y}^{0}=-1$). Simulations are for (a) $\varepsilon=1$, (b) $\varepsilon=0.1$, and (c) $\varepsilon=0.01$. Top panels display the dynamics of the rescaled phenotype density $n(t,y)/\rho(t)$; bottom panels display the phenotype density $n(t,y)$ (red solid line) at select times. We compare the numerical solution with the analytically predicted one (black dashed lines) – i.e. the Gaussian-like solution (38) with mean $\bar{y}_{\varepsilon}(t)$ and weight $\rho_{\varepsilon}(t)$ obtained by solving numerically the Cauchy problem (39). We see the solution to be of a Gaussian form, which becomes concentrated as a weighted increasingly sharp Gaussian as $\varepsilon$ gets smaller, corroborating the emergence of concentration phenomena of type (40). We see also that the mean or prevailing phenotype $\bar{y}(t)$ converges to $y=\varphi$ (black dashed lines, top panels) as $t\to\infty$, corresponding to selection of the fittest trait.
@@
Finally, for a spatially heterogeneous environment – for example a population surrounded by non-uniformly distributed abiotic factors (nutrients, therapeutic agents, etc) that affects proliferation and death rates – a natural extension of PS-PDE models of form (7) is